OpenStudy (anonymous):

\displaystyle\lim_{x\rightarrow\infty} \frac{9x}{x-2}

OpenStudy (anonymous):

if you divide (long division) you'll see that you'll get 9.

OpenStudy (anonymous):

\[\lim_{x\rightarrow\infty}\frac{9x}{x-2}=9\lim_{x\rightarrow\infty}\frac{x}{x-2}=9(1)=9\]

jimthompson5910 (jim_thompson5910):

or you can divide each term by x to go from \[\large \frac{9x}{x-2}\] to \[\large \frac{9}{1-\frac{2}{x}}\]

jimthompson5910 (jim_thompson5910):

then notice as x ---> infinity, then 2/x ---> 0 so that's why the limiting value is 9

OpenStudy (anonymous):

\[\frac{9x}{x-2}=9+\frac{18}{x-2}\] \[\Rightarrow \lim_{x \rightarrow \infty}\frac{ 9x }{ x-2 }=\lim_{x \rightarrow \infty}\left[9+\frac{ 18 }{ x-2 }\right]=9+0=9\]